OpenStudy (anonymous):

The perimeter of a triangle is 240. Its sides are proportional to 3:4:5. What is the largest side of the triangle?

OpenStudy (mathmagician):

3x+4x+5x=240, therefore x=20. So, the longest side is 5x=5*20=100

OpenStudy (anonymous):

right